If the domain of f

f is -3\le x \le 3

−3≤x≤3, what is the range


1
Expert's answer
2022-05-06T13:03:24-0400

Let "f(x)=x^2" and domain of f is "-3\le x\le 3."

Then the range is "0\le y\le 9."